\begin{thebibliography}{99}
\bibitem{Akhiezer_Krein1937}  Akhiezer N.I., Krein M. G.
\emph{On the best approximation of differentiable periodic functions by trigonometric
	polynomials.} Dokl. Akad. Nauk SSSR 1937, \textbf{15} (3), 107-112. (in Russian)

\bibitem{Bushanskij} Bushanskii A.V. \emph{Best harmonic approximation in the mean of certain functions.} Studies in the theory of approximation of functions and their applications, Akad. Nauk Ukrain. SSR, Inst. Mat., Kiev, 1978, 29-37. (in Russian)

\bibitem{Dzyadyk _1959}   Dzyadyk V.K. \emph{On the best approximation on the classes of periodic functions determined by kernels which are integrals of absolutely monotone functions.} Izv. Akad. Nauk SSSR, Ser. Mat. 1959, \textbf{23} (6), 933-950. (in Russian)

\bibitem{Dzyadyk _1974}   Dzyadyk V.K. \emph{Best approximation on classes of periodic functions defined by integrals of linear combinations of absolutely monotonous kernels.} Mat. Zametki 1974, \textbf{16} (5), 691-701. 

\bibitem{Hrabova_Serdyuk2013} Hrabova U.Z.,   Serdyuk A.S. \emph{Order estimates for the best approximations and approximations by Fourier sums of the classes of $(\psi,\beta)$-differential functions}. Ukrainian Math. J. 2014, \textbf{65} (9), 1319-1331. https://doi.org/10.1007/s11253-014-0861-7

\bibitem{Favard_1936} Favard J.  \emph{Sur l'approximation
	des fonctions p$\acute e$riodiques par des polynomes trigonom$\acute e$triques.} C.R.Acad. Sci. 1936, \textbf{203}, 1122-1124. (in French)

\bibitem{Favard_1937} Favard J.  \emph{Sur les meilleurs
	proc$\acute e$des d'approximations de certains classes de fontions
	par des polynomes trigonometriques.} Bull. de Sci. Math. 1937, \textbf{61}, 209--224, 243 -- 256. (in French)

\bibitem{Kornejchuk1987}   Korneichuk N.P. {Exact constants in approximation theory.} Encyclopedia of Mathematics and its Applications, 38. Cambridge University Press, Cambridge, 1991. 

\bibitem{Kornejchuk1992}  Korneichuk N.P.,  Babenko V.F.,  Ligun A.A.  {Extremal properties of polynomials and splines.} Naukova Dumka, Kiev, 1992. (in Russian)

\bibitem{Krein_1938} Krein  M.G. \emph{The theory of best approximation of periodic functions.} Dokl. Akad. Nauk SSSR 1938, \textbf{18} (4-5), 245-249. (in Russian)

\bibitem{Kushpel'1989_UMZh4}  Kushpel' A.K.  \emph{Estimates for the widths of classes of analytic functions.} Ukrainian Math. J. 1989, \textbf{41} (4), 493-496. https://doi.org/10.1007/BF01060633


\bibitem{Sz.-Nagy1938}
Sz.-Nagy B. 
\emph{\"Uber gewisse Extremalfragen bei transformierten trigonometrischen Entwicklungen. 1. Periodischer Fall},
Ber. Math.-Phys. Kl. Akad. Wiss., Leipzig, 1938, \textbf{90}, 103--134,
(in German)

\bibitem{Pinkus1985}  Pinkus A. {n-widths in approximation theory}, Springer-Verlag, Berlin, 1985. 

\bibitem{Romanyuk2012} Romanyuk A.S. Approximating Characteristics of the Classes of Periodic Functions of Many Variables. Proceedings of the Institute of Mathematics of the National Academy of Sciences of Ukraine. Kyiv, 2012. (in Russian)

\bibitem{Serdyuk1995}  Serdyuk A.S. \emph{On the best approximation of classes of convolutions of periodic functions by trigonometric polynomials.} Ukrainian Math. J. 1995, \textbf{47} (9), 1435-1440. https://doi.org/10.1007/BF01057518

\bibitem{Serdyuk1998} Serdyuk A.S. \emph{Estimates for the widths and best approximations of classes of convolutions of periodic functions.} Fourier series: theory and applications  (Kamenets-Podol'skii, 1997), Pr. Inst. Mat. Nats. Akad. Nauk Ukr. Mat. Zastos. 1998, \textbf{20}, 286-299. (in Ukrainian)

\bibitem{Serdyuk1999}  Serdyuk A.S. \emph{Widths and best approximations for classes of convolutions of periodic functions.} Ukrainian Math. J. 51, no. 5  (1999), 748-763. https://doi.org/10.1007/BF02591709

\bibitem{Serdyuk_2002_zb} Serdyuk A.S. \emph{On best approximation in classes of convolutions of periodic functions.} Theory of the approximation of functions and related problems. Pr. Inst. Mat. Nats. Akad. Nauk Ukr. Mat. Zastos. 2002, \textbf{35}, 172-194. (in Ukrainian)

\bibitem{Serdyuk_2005_7} Serdyuk A.S. \emph{Best approximations and widths of classes of convolutions of periodic functions of high smoothness.} Ukrainian Math. J. 2005, \textbf{57} (7), 1120-1148. https://doi.org/10.1007/s11253-005-0251-2 

\bibitem{Serdyuk_2005_8} Serdyuk A.S. \emph{Approximation of classes of analytic functions by Fourier sums in the uniform metric.} Ukrainian Math. J. 2005, \textbf{57} (8),  1275-1296. https://doi.org/10.1007/s11253-005-0261-0

\bibitem{Serdyuk_Bodenchuk2013JAT} Serdyuk A.S.,  Bodenchuk V.V. \emph{Exact values of Kolmogorov widths of classes of Poisson integrals}. J. Approx. Theory 2013, \textbf{173}, 89-109. https://doi.org/10.1016/j.jat.2013.05.002

\bibitem{Serdyuk_Sokolenko2011JAT} Serdyuk A.S.,  Sokolenko I.V. \emph{Asymptotic behavior of best approximations of classes of Poisson integrals of functions from $H_\omega$.} J. Approx. Theory 2011, \textbf{163}, 1692-1706. https://doi.org/10.1016/j.jat.2011.06.008

\bibitem{Serdyuk_Sokolenko2011Zb} Serdyuk A.S. Sokolenko I.V.
\emph{Uniform approximation  of classes of $(\psi,\bar\beta)-$differentiable functions by linear methods.}  {Approximation Theory of Functions and Related Problems, Zb. prats Inst. mat. NAN Ukr., Kyiv} 2011, \textbf{8}, (1),	181--189. (in Ukrainian)

\bibitem{Serdyuk_Sokolenko2020}	Serdyuk A.S., Sokolenko I.V. \emph{Asymptotic estimates for the best uniform approximations of 	classes of convolution of periodic functions of high smoothness.}  Journal of Mathematical Sciences  2021, \textbf{252} (4), 526-540. https://doi.org/10.1007/s10958-020-05178-1

\bibitem{Serdyuk_Stepanyuk2014_UMZh9} Serdyuk A.S.,  Stepanyuk T.A.  \emph{Estimations of the best approximations for the classes of infinitely differentiable functions in uniform and integral metrics.} Ukrainian Math. J. 2015,  \textbf{66} (9), 1393-1407. https://doi.org/10.1007/s11253-015-1134-9

\bibitem{Serdyuk_Stepanyuk2014_UMZh12} Serdyuk A.S.,  Stepanyuk T.A.  \emph{Order estimates for the best approximations and approximations by Fourier sums in the classes of convolutions of periodic functions of low smoothness in the uniform metric.} Ukrainian Math. J. 2015, \textbf{66} (12), 1862-1882. https://doi.org/10.1007/s11253-015-1056-6

\bibitem{Serdyuk_Stepanyuk2019} Serdyuk A.S.,  Stepanyuk T.A. \emph{Uniform Approximations by Fourier Sums in Classes of Generalized Poisson Integrals.} Analysis Math. 2019, \textbf{45} (1), 201-236. https://doi.org/10.1007/s10476-018-0310-1

\bibitem{Shevaldin_1992}  Shevaldin V.T. \emph{Widths of classes of convolutions with Poisson kernel.}  Math. Notes 1992, \textbf{51} (6), 611-617. https://doi.org/10.1007/BF01263308

\bibitem{Stechkin1956_2}  Stechkin S.B. \emph{On the best approximation of certain classes of periodic functions by trigonometric polynomials}, Izv. Akad. Nauk SSSR, Ser. Mat. 1956, \textbf{20}, 643--648. (in Russian)

\bibitem{Stepanets1987} Stepanets A.I. {Classification and Approximation of Periodic Functions.} Kluwer Academic Publishers, Dordrecht, 1995.		

\bibitem{Stepanets2005} Stepanets A.I. {Methods of Approximation Theory.} Utrecht, VSP,
2005.

\bibitem{Stepanets_Serdyuk1995UMZh8} Stepanets A.I., Serdyuk A.S. \emph{Lower bounds for the widths of classes of convolutions of periodic functions in $C$ and $L$ metrics}. Ukrainian Math. J. 1995, \textbf{47} (8), 1271-1282. https://doi.org/10.1007/BF01057715 

\bibitem{Sun_1961} Sun Yong-sheng, \emph{On the best approximation of periodic differentiable functions by trigonometric polynomials. II.} Izv. Akad. Nauk SSSR Ser. Mat. 1961, \textbf{25} (1),  143-152. (in Russian)


\bibitem{Temlyakov1990} Temlyakov V.N.  \emph{On estimates for the widths of classes of infinitely differentiable functions.}  Mat. Zametki 1990, \textbf{47} (5),  155-157. (in Russian)	

\bibitem{Temlyakov1993} Temlyakov V.N. {Approximation of periodic functions}, Comput. Math. Anal. Ser., Nova Science Publishers, Inc., Commack, NY, 1993.

\bibitem{Tikhomirov1976} Tikhomirov V.M.  \emph{Some questions in approximation theory}. Some questions in approximation theory. Izdat. Moskov. Univ., Moscow, 1976. (in Russian)	

\end{thebibliography}

\documentclass[draft,11pt,english]{article}

\usepackage{amsfonts,amsmath,amsthm,amscd,amssymb,latexsym,cite,verbatim,texdraw,floatflt,
caption2,pb-diagram}
\usepackage[mathscr]{eucal}

\theoremstyle{plain}
\newtheorem{theorem}{Theorem}%[section]

\newtheorem{corollary}{Corollary}[section]

\theoremstyle{definition}

\newcommand{\keywords}{\textbf{Key words. }\medskip}
\newcommand{\subjclass}{\textbf{MSC 2010. }\medskip}
\renewcommand{\abstract}{\textbf{Abstract. }\medskip}

\begin{document}

\sloppy

\title{Asymptotic estimates for the widths \\ of classes of functions of high smoothness \thanks{This work was partially supported by the Grant H2020-MSCA-RISE-2019, project number 873071 (SOMPATY: Spectral Optimization: From Mathe\-ma\-tics to Physics and Advanced Technology).%Authors want to thanks \dots
}}

\author{A.S.~Serdyuk and I.V.~Sokolenko}

%\shorttitle{Asymptotic estimates for best uniform approximations}

%\shortauthor{A.S.~Serdyuk, I.V.~Sokolenko}

\date{\ }

\maketitle

\begin{abstract}
We find two-sided estimates for Kolmogorov, Bernstein, linear and projection widths of the classes of  convolutions of $2\pi$-periodic functions $\varphi$, such that $\|\varphi\|_2\le1$, with fixed generated kernels $\Psi_{\bar{\beta}}$, which have Fourier series of the form $\sum\limits_{k=1}^\infty \psi(k)\cos(kt-\beta_k\pi/2), $ where  $\psi(k)\ge0,$ $\sum\psi^2(k)<\infty, \beta_k\in\mathbb{R},$ in the space $C$.
It is shown that for  rapidly decrising sequences $\psi(k)$ (in particular, if $\lim\limits_{k\rightarrow\infty}{\psi(k+1)}/{\psi(k)}=0$) obtained estimates are asymptotic equalities. We establish that asymptotic equalities for  widths of this classes are realized by trigonometric Fourier sums.
\end{abstract}

\subjclass{41A46, 42A10.}

\keywords{Bernstein width, Kolmogorov width, linear width, projection width, Fourier sum, Weyl-Nagy class, class of the generalized Poisson integrals, $(\psi,\bar{\beta})$-integral, asymptotic equality}

\section*{Introduction}
Let   $L_{p}$,
$1\le  p<\infty$, be the space of $2\pi$--periodic functions $f$ sum\-mable to the power $p$ on  $[-\pi,\pi)$, in which the norm is given by the formula
$$
\|f\|_{L_p}=\|f\|_{p}=\bigg(\int\limits_{-\pi}^{\pi}|f(t)|^pdt\bigg)^{1/p},
$$
$L_{\infty}$ be the space of measurable and essentially bounded   $2\pi$--periodic functions  $f$ with the norm
$$
\|f\|_{L_\infty}=\|f\|_{\infty}=\mathop {\rm ess \sup}\limits_{t} |f(t)|,
$$
$C$ be the space of continuous $2\pi$--periodic functions  $f$, in which the norm is defined by the equality
$$
{\|f\|_{C}=\max\limits_{t}|f(t)|}.
$$

Denote by $C^\psi_{\bar\beta,p},\ 1\le p\le\infty,$ the set of all $2\pi$-periodic functions $f$,
representable as convolution
\begin{equation}\label{25_7'}
	f(x)=\frac{a_0}{2}+\frac{1}{\pi}\int\limits_{-\pi}^{\pi}
	\varphi(x-t) \Psi_{\bar\beta}(t)dt, \ \ \ a_0\in\mathbb R, \ \ \ \varphi\in B_p^0,
\end{equation}
$$
B_p^0=\{g\in L_p:\ \|g\|_p\le1,\ g\perp1\},
$$
with a fixed generated kernel $\Psi_{\bar\beta}\in L_{p'},\  1/p+ 1/{p'}=1,\ $ the Fourier series of which has the form
\begin{equation}\label{1*}
	S[\Psi_{\bar{\beta}}](t)=\sum\limits_{k=1}^\infty \psi(k)\cos\left(kt-\frac{\beta_k\pi}2\right),\quad \beta_k\in\mathbb{R},\quad \psi(k)\ge0.
\end{equation}
A function $f$ in the representation (\ref{25_7'}) is called $(\psi,\bar{\beta})$-integral of the function $\varphi$ and  is denoted by ${\mathcal J}^{ \psi}_{\bar{\beta}}\varphi$  $(f={\mathcal J}^{ \psi}_{\bar{\beta}}\varphi)$. If $\psi(k)\neq0,\ k\in\mathbb{N},$ then the function $\varphi$ in the representation  (\ref{25_7'}) is called $(\psi,\bar{\beta})$-derivative of the function $f$ and is denoted by $f^{ \psi}_{\bar{\beta}}$ $(\varphi=f^{ \psi}_{\bar{\beta}})$. The concepts of  $(\psi,\bar{\beta})$-integral and $(\psi,\bar{\beta})$-derivative  was introduced by Stepanets (see, e.g., \cite{Stepanets1987,Stepanets2005}).
Since \mbox{$\varphi\in L_p$} and $\ \Psi_{\bar\beta}\in L_{p'},$ then  the function $f$ of the form \eqref{25_7'} is a continuous function, i.e. $C^\psi_{\bar\beta,p}\subset C$ (see \cite[Proposition 3.9.2.]{Stepanets2005}).

In the case $\beta_k\equiv\beta,\ \beta\in \mathbb R,$ the classes $C^\psi_{\bar\beta,p}$ are denoted by $C^\psi_{\beta,p}$.

For $\psi(k)=k^{-r}, r>0, $ the classes  $C^\psi_{\bar\beta,p}$ and $C^\psi_{\beta,p}$ are denoted by $W^r_{\bar\beta,p}$ and $W^r_{\beta,p}$, respectively. The classes $W^r_{\beta,p}$ are  the well-known  Weyl-Nagy classes (see, e.g., \cite{Sz.-Nagy1938, Stechkin1956_2, Stepanets1987, Stepanets2005}). In other words, $W^r_{\beta,p}, 1\le p\le \infty,$ are the classes of  $2\pi$-periodic functions $f$, representable
as convolutions of the form
\begin{equation}\label{1}
	f(x)=\frac{a_0}{2}+\frac{1}{\pi}\int\limits_{-\pi}^{\pi}
	\varphi(x-t) B_{r,\beta}(t)dt, \ \ \ a_0\in\mathbb R,
\end{equation}
the Weyl-Nagy kernels $B_{r,\beta}$ of the form
\begin{equation}\label{2*}
	B_{r,\beta}(t)=\sum\limits_{k=1}^\infty k^{-r}\cos\left(kt-\frac{\beta\pi}2\right),\quad r>0,\quad \beta\in\mathbb{R},
\end{equation}
with functions $\varphi \in B_p^0$. The function $\varphi$ in the formula \eqref{1} is called  the Weyl-Nagy derivative of the function $f$ and is denoted by $f_\beta^r$.

If  $r\in\mathbb N$ and $ \beta=r,\ $ then the functions $B_{r,\beta}$ of the form (\ref{2*})  are  the well-known Bernoulli kernels and  the corresponding classes $W^r_{\beta,p}$ coincide with the well-known classes $W^r_{p}$ which consist of $2\pi$-periodic functions $f$  with absolutely continuous derivatives $f^{(k)}$ up to $(r-1)$-th order inclusive and such that  $\|f^{(r)}\|_p\le1$. In addition,   for almost everywhere $x\in\mathbb{R} \ \ f^{(r)}(x)=f_r^r(x)=\varphi(x),\ $  where $\varphi$ is the function from (\ref{1}).

For $\psi(k)=e^{-\alpha k^{r}},\ \alpha>0, \ r>0, $   the classes  $C^\psi_{\bar\beta,p}$ and $C^\psi_{\beta,p}$ are denoted by $C^{\alpha,r}_{\bar\beta,p}$ and $C^{\alpha,r}_{\beta,p}$, respectively.  The sets $C^{\alpha,r}_{\beta,p}$ are  well-known classes of the generalized Poisson integrals (see, e.g., \cite{Stepanets1987, Stepanets2005}), i.e. classes of  convolutions
\begin{equation}\label{'1}
	f(x)=\frac{a_0}{2}+\frac{1}{\pi}\int\limits_{-\pi}^{\pi}
	\varphi(x-t) P_{\alpha,r,\beta}(t)dt, \quad a_0\in\mathbb R,\quad \varphi \in B_p^0,
\end{equation}
with the generalized Poisson kernels
\begin{equation}\label{'2*}
	P_{\alpha,r,\beta}(t)=\sum\limits_{k=1}^\infty e^{-\alpha k^{r}} \cos\left(kt-\frac{\beta\pi}2\right),\quad \alpha>0, \quad r>0,\quad \beta\in\mathbb{R}.
\end{equation}

Let $\mathfrak N$ be a some functional class from the space $C$ $(\mathfrak N\subset C)$. The quantity
\begin{equation}\label{15_3'}
	E_n(\mathfrak N)_C=\sup\limits_{f\in \mathfrak N} E_n(f)_C= \sup\limits_{f\in \mathfrak N}\inf\limits_{T_{n-1}\in{\mathcal T}_{2n-1}} \|f-T_{n-1}\|_C
\end{equation}
is called the best uniform approximation of the class $\mathfrak{N}$ by elements of the
subspace ${\mathcal T}_{2n-1}$ of trigonometric polynomials $T_{n-1}$ of the order $n-1$:
$$
T_{n-1}(x)=\frac{\alpha_0}{2}+\sum\limits_{k=1}^{n-1}
(\alpha_k\cos kx+\beta_k\sin kx),\ \ \  \alpha_k,\beta_k\in \mathbb{R}.
$$

The order estimates for the best approximations $E_n(K)_C$ of classes \mbox{$K=C^\psi_{\bar\beta,p},$} $ 1\le p\le\infty,$ (and, hence, classes $W^r_{\beta,p}$, $C^{\alpha,r}_{\beta,p}$ and $C^\psi_{\beta,p}$) depending on rate of decreasing to zero of sequences $\psi(k)$  were obtained, in particular, in the works of Temlyakov \cite{Temlyakov1993}, Hrabova and Serdyuk \cite{Hrabova_Serdyuk2013}, Serdyuk and Stepanyuk \cite{Serdyuk_Stepanyuk2014_UMZh9,Serdyuk_Stepanyuk2014_UMZh12}.% in the works \cite{Temlyakov1993,Hrabova_Serdyuk2013,Serdyuk_Stepanyuk2014_UMZh9,Serdyuk_Stepanyuk2014_UMZh12}.

If the sequences $\psi(k)$ decrease to zero faster than any geometric progression, then asymptotic equations of the best uniform approximations are even known (see, for example, the authors work \cite{Serdyuk_Sokolenko2020} and the bibliography available there).

In \cite{Serdyuk_Sokolenko2020} it was shown that for such classes $C^\psi_{\bar\beta,p}$ the following asymptotic equations take places
\begin{equation}\label{**}
	E_n(C^\psi_{\bar\beta,p})_C\sim {\mathcal E}_n(C^\psi_{\bar\beta,p})_C\sim\frac{\|\cos t\|_{p'}}{\pi}\psi(n), \quad1\le p\le\infty,	
\end{equation}
where
\begin{equation*}\label{key}
	{\mathcal E}_n(C^\psi_{\bar\beta,p})=\sup_{f\in C^\psi_{\bar\beta,p}}\|f-S_{n-1}(f)\|_C,
\end{equation*}
$S_{n{-}1}(f)$ is the partial Fourier sum of order $n{-}1$ of the function $f$, \mbox{$\displaystyle\frac1p+\frac1{p'}=1,$} and %"$\sim$"\ is asymptotically equivalent to  i.e.
$A(n){\sim} B(n)$ as $n\rightarrow\infty$ means that $\lim\limits_{n\rightarrow\infty} {A(n)}/{B(n)}{=}1.$

For $p=\infty$ in the case of $K=W^r_{\bar\beta,\infty}, r>0, $ and in the cases of $K=C^{\alpha,r}_{\bar\beta,\infty}, r\ge1,$ and  $K=C^\psi_{\bar\beta,\infty}$ $(K=C^\psi_{\beta,\infty})$  for certain restrictions on sequences $\psi$ and $\bar{\beta}$ the exact values of the best uniform approximations are known thanks to the works  of Favard
\cite{Favard_1936,Favard_1937},   Akhiezer and Krein \cite{Akhiezer_Krein1937}, Krein \cite{Krein_1938}, Nagy \cite{Sz.-Nagy1938}, Stechkin \cite{Stechkin1956_2}, Dzyadyk   \cite{Dzyadyk _1959,Dzyadyk _1974}, Sun \cite{Sun_1961}, Bushanskij \cite{Bushanskij}, Pinkus \cite{Pinkus1985}, Serdyuk \cite{Serdyuk1995,Serdyuk1998,Serdyuk1999,Serdyuk_2002_zb,Serdyuk_2005_7}    etc.%\cite{Favard_1936,Favard_1937,Sz.-Nagy1938,Akhiezer_Krein1937,Dzyadyk _1959,Dzyadyk _1974,Krein_1938,Stechkin1956_2,Sun_1961,Bushanskij,Serdyuk1995,Serdyuk1999,Serdyuk_2002_zb,Pinkus1985}.

For $p=2$ and for arbitrary $\bar{\beta}=\beta_k\in\mathbb{R}, \sum\limits_{k=1}^\infty\psi^2(k)<\infty$ the exact values for the quantity ${\mathcal E}_n(C^\psi_{\bar\beta,2})_C$ are also known  (see  \cite{Serdyuk_Sokolenko2011Zb}).

In this paper, we establish two-sided estimates of Kolmogorov, Bernstein, linear and projection widths of the classes $C^\psi_{\bar\beta,2}$ in the space $C$, which become into asymptotic equations under certain restrictions on the sequence $\psi(k)$ (in particular, if $\lim\limits_{k\rightarrow\infty}{\psi(k+1)}/{\psi(k)}=0$).

Let  $K$ be a convex centrally symmetric subset of $C$ and
let ${ B }$ be a unit ball of the space  $C$. Let also  $F_N$ be an arbitrary $N$-dimensional subspace of space $C$, $N\in {\mathbb N}$, and
$\mathscr{L}(C, F_N)$ be a set of linear operators from $C$ to $F_N$.
By    $\mathscr {P}(C, F_N)$ denote the subset of projection operators of the set ${\mathscr{L}}(C, F_N)$,  that is, the set of the operators  $A$ of linear projection onto the set $F_N$ such that $Af = f$ when $f\in F_N$.  The quantities
\begin{equation}\label{A}
	b_N(K, C)=\sup\limits _{F_{N+1}}\sup\{\varepsilon>0: \varepsilon { B }\cap F_{N+1}
	\subset K\},
\end{equation}
\begin{equation}\label{B}
	d_N(K, C)=\inf\limits _{F_N}\sup \limits _{f\in K}
	\inf \limits _{u\in F_N}\|f - u \|_{_{\scriptstyle  C}} ,
\end{equation}
\begin{equation}\label{C}
	\lambda _N(K,C)=
	\inf \limits _{F_N}\inf \limits_{A\in {\mathscr {L}}(C, F_N)}\sup \limits _{f\in K}
	\|f - Af\|_{_{\scriptstyle  C}} ,
\end{equation}
\begin{equation}\label{D}
	\pi _N(K, C)=\inf \limits_{F_N}\inf \limits _{A\in {\mathscr {P}}(C,F_N)}
	\sup \limits _{f\in K}\|f - Af\|_{_{\scriptstyle  C}} ,
\end{equation}
are called Bernstein, Kolmogorov, linear, and projection $N$-widths of the set $K$ in the space $C$, respectively.

The results containing order estimates of the widths \eqref{A}-\eqref{D} in the case of $K=C_{\bar\beta,p}^\psi$ (and, in particular, $W^r_{\beta,p}$ and $C_{\beta,p}^\psi$) can be found, for example, in the works  of Tikhomirov \cite{Tikhomirov1976}, Pinkus \cite{Pinkus1985}, Kornejchuk \cite{Kornejchuk1987}, Kushpel' \cite{Kushpel'1989_UMZh4}, Romanyuk \cite{Romanyuk2012}, Temlyakov \cite{Temlyakov1990,Temlyakov1993}  etc. 

\section{Main results}
The main result of this paper is the following statement.

\begin{theorem}\label{theorem1}
	Let $\bar{\beta}=\{\beta_k\}_{k=1}^\infty,$ $\beta_k\in\mathbb R,$ and  $\psi(k)>0$  satisfies the condition
	\begin{equation}\label{1t1}
		\sum\limits_{k=1}^\infty\psi^2(k)<\infty.
	\end{equation}
	Then for all $n\in\mathbb{N}$ the following inequalities hold
	\begin{equation}
		\frac1{\sqrt{\pi}}\left(
		\frac1{\psi^2(n)}+2\sum\limits_{k=1}^{n-1}\frac1{\psi^2(k)}\right)^{-\frac12}\le P_{2n}(C_{\bar\beta,2}^\psi,C)
		\le P_{2n-1}(C_{\bar\beta,2}^\psi,C)%\le
		\le \frac1{\sqrt{\pi}}\left(\sum\limits_{k=n}^\infty\psi^2(k)\right)^{\frac12},
		\label{1t2}
	\end{equation}
	where $P_N$ is any of the widths $b_N, d_N, \lambda_N$ or $\pi_N$.
	\\
	\indent If, in adition, $\psi(k)$ satisfies the condition
	\begin{equation}\label{1t3}
		\lim\limits_{n\rightarrow\infty}\max\left\{
		\psi(n)\left(\sum\limits_{k=1}^{n-1}\frac1{\psi^2(k)}\right)^{\frac12},
		\frac1{\psi(n)}{\left(\sum\limits_{k=n+1}^\infty\psi^2(k)\right)^{\frac12}}	\right\}=0,
	\end{equation}
	then the following asymptotic equalities hold
	\begin{equation}\label{1t4}
		\begin{split}
			\left.\begin{array}{l}
				P_{2n}(C_{\bar\beta,2}^\psi,C)\\
				P_{2n-1}(C_{\bar\beta,2}^\psi,C)
			\end{array}\right\}= & \psi(n)\Bigg(\frac1{\sqrt{\pi}}
			\\
			+& 		\left.{\mathcal O}(1) \max\left\{	\psi(n)\left(\sum\limits_{k=1}^{n-1}\frac1{\psi^2(k)}\right)^{\frac12},
			\frac1{\psi(n)}{\left(\sum\limits_{k=n+1}^\infty\psi^2(k)\right)^{\frac12}}\right\} \right),
		\end{split}
	\end{equation}
	where ${\mathcal O}(1)$ are the quantities uniformly bounded in all parameters.
	\\
	\indent The equalities \eqref{1t4} are realized by trigonometric Fourier sums $S_{n-1}(f)$.
\end{theorem}

\begin{proof} In the work \cite{Serdyuk_Sokolenko2011Zb} it was proved that if the condition \eqref{1t1} is satisfied then the following equality holds
	\begin{equation}\label{1td1}
		{\mathcal E}(C_{\bar\beta,2}^\psi;S_{n-1})_C=\frac1{\sqrt{\pi}}	\left(\sum\limits_{k=n}^\infty\psi^2(k)\right)^{\frac12},\quad\beta_k\in\mathbb{R},\quad n\in\mathbb{N}.
	\end{equation}

	Since the operator that assigns to each function $f\in C$ its partial Fourier sum is a linear projector, then by virtue \eqref{1td1}
	\begin{equation}\label{1td2}
		\pi_{2n-1}(C_{\bar\beta,2}^\psi,C)\le {\mathcal E}(C_{\bar\beta,2}^\psi;S_{n-1})_C=\frac1{\sqrt{\pi}}	\left(\sum\limits_{k=n}^\infty\psi^2(k)\right)^{\frac12},\quad\beta_k\in\mathbb{R},\quad n\in\mathbb{N}.
	\end{equation}
	
	For all  $n\in\mathbb{N}$ and $\mathfrak{N}\subset C$
	\begin{equation*}\label{key}
		P_{2n}(\mathfrak{N},C)\le
		P_{2n-1}(\mathfrak{N},C),
	\end{equation*}
	where $P_N$ is any of the widths $b_N, d_N, \lambda_N,$ and $\pi_N$, and, in addition, for all $N\in\mathbb{N}$
	\begin{equation}\label{1td3}
		b_N(\mathfrak{N}, C)\le d_N(\mathfrak{N}, C)\le\lambda_N(\mathfrak{N}, C)\le\pi_N(\mathfrak{N}, C).
	\end{equation}
	Therefore on the basis of \eqref{1td2} we obtaine an estimate from above for the widths $P_N$ in the formula  \eqref{1t2}.
	
	To obtain a required estimate from below in \eqref{1t2}  it suffices to establish that
	\begin{equation}\label{1td4}
		b_{2n}(C_{\bar\beta,2}^\psi,C)\ge \frac1{\sqrt{\pi}}\left(
		\frac1{\psi^2(n)}+2\sum\limits_{k=1}^{n-1}\frac1{\psi^2(k)}\right)^{-\frac12}.
	\end{equation}
	
	In $(2n+1)$-dimensional space ${\mathcal T}_{2n+1}$ of trigonometric polynomials $ T_n $ of order $ n $ let us consider a ball of the form
	\begin{equation}\label{1td5}
		B_{2n+1}=\left\{T_n\in{\mathcal T}_{2n+1}:\|T_n\|_C\le \frac1{\sqrt{\pi}}\left(
		\frac1{\psi^2(n)}+2\sum\limits_{k=1}^{n-1}\frac1{\psi^2(k)}\right)^{-\frac12}\right\}
	\end{equation}
	and prove the following embedding
	\begin{equation}\label{1td6}
		B_{2n+1}\subset C_{\bar\beta,2}^\psi.
	\end{equation}
	
	For any trinometric polynomial
	\begin{equation}\label{1td7}
		T_n(x)=\frac{a_0}2+\sum\limits_{k=1}^n (a_k\cos kx+b_k\sin kx)
	\end{equation}
	from the ball $B_{2n+1}$ its $(\psi,\bar{\beta})$-derivative has a form
	$$
	(T_n)_{\bar{\beta}}^\psi(x)=\sum\limits_{k=1}^n \left(\frac{a_k}{\psi(k)}\cos \left(kx+\frac{\beta_k\pi}{2}\right)+\frac{b_k}{\psi(k)}\sin \left(kx+\frac{\beta_k\pi}{2}\right)\right)
	$$
	$$
	=\sum\limits_{k=1}^n \left(\frac{a_k\cos \frac{\beta_k\pi}{2}}{\psi(k)}\cos kx-
	\frac{a_k\sin \frac{\beta_k\pi}{2}}{\psi(k)}\sin kx+
	\frac{b_k\cos\frac{\beta_k\pi}{2}}{\psi(k)}\sin kx+
	\frac{b_k\sin\frac{\beta_k\pi}{2}}{\psi(k)}\cos kx\right)
	$$
	\begin{equation}\label{1td8}
		=\sum\limits_{k=1}^n \frac1{\psi(k)}\left(\left({a_k\cos \frac{\beta_k\pi}{2}}
		+{b_k\sin\frac{\beta_k\pi}{2}}\right)\cos kx+\left(-
		{a_k\sin \frac{\beta_k\pi}{2}}+{b_k\cos\frac{\beta_k\pi}{2}}\right)\sin kx\right).
	\end{equation}
	
	By virtue of Parseval equality, from \eqref{1td8} we get
	\begin{equation}\label{1td9}
		\left\|(T_n)_{\bar{\beta}}^\psi\right\|_2
		{=}\sqrt{\pi}	\left(\sum\limits_{k=1}^n \frac1{\psi^2(k)}\left(\left({a_k\cos \frac{\beta_k\pi}{2}}
		{+}{b_k\sin\frac{\beta_k\pi}{2}}\right)^2{+}\left({-}
		{a_k\sin \frac{\beta_k\pi}{2}}{+}{b_k\cos\frac{\beta_k\pi}{2}}\right)^2\right)\right)^{\frac12}
		$$
		$$
		=\sqrt{\pi}	\left(\sum\limits_{k=1}^n \frac1{\psi^2(k)}\left(a_k^2+b_k^2\right)\right)^{\frac12}.
	\end{equation}
	
	By Parseval equality for the polynomial $ T_n $ of the form \eqref{1td7} we obtain
	\begin{equation}\label{1td10}
		\frac{a_0^2}2+\sum\limits_{k=1}^n (a_k^2+b_k^2)=\frac1\pi\int\limits_{-\pi}^{\pi}T_n^2(x) dx.
	\end{equation}
	Therefore we have a chain of inequalities
	\begin{equation}\label{1td11}
		a_k^2+b_k^2\le \sum\limits_{k=1}^n (a_k^2+b_k^2)\le \frac1\pi\int\limits_{-\pi}^{\pi}\left\|T_n\right\|_C^2 dx= 2\left\|T_n\right\|_C^2,
	\end{equation}
	and, consequently, we obtaine an estimate for $\sqrt{a_k^2+b_k^2}$ of the following form
	\begin{equation}\label{1td12}
		\sqrt{a_k^2+b_k^2}\le\sqrt2\left\|T_n\right\|_C,\quad k=\overline{1,n}.
	\end{equation}
	
	In the case of $ k = n $ this estimate  can be improved. To do this, let us consider a trigonometric polynomial
	$$
	\tau_n(x):=\frac{T_n(x)}{\sqrt{a_n^2+b_n^2}}=\frac1{\sqrt{a_n^2+b_n^2}}\left(\frac{a_0}2+\sum\limits_{k=1}^n (a_k\cos kx+b_k\sin kx)\right)
	$$
	$$
	=\frac1{\sqrt{a_n^2+b_n^2}}\left(\frac{a_0}2+\sum\limits_{k=1}^n \sqrt{a_k^2+b_k^2} (\frac{a_k}{\sqrt{a_k^2+b_k^2}}\cos kx+\frac{b_k}{\sqrt{a_k^2+b_k^2}}\sin kx)\right)
	$$
	\begin{equation}\label{1td13}
		=\frac{\rho_0}{2}+\sum\limits_{k=1}^n\rho_k\cos(kx+\theta_k),
	\end{equation}
	where
	$$
	\rho_0=\frac{a_0}{\sqrt{a_n^2+b_n^2}},\quad\rho_k=\frac{\sqrt{a_k^2+b_k^2}}{\sqrt{a_n^2+b_n^2}},\quad k=\overline{1,n-1},\quad\rho_n=1,
	$$
	and $\theta_k$ are such that
	$$
	\left\{\begin{array}{rr}
		\cos\theta_k=\displaystyle\frac{a_k}{\sqrt{a_k^2+b_k^2}}, \ k=\overline{1,n},
		\\
		\sin\theta_k=\displaystyle\frac{-b_k}{\sqrt{a_k^2+b_k^2}}, \ k=\overline{1,n}.
	\end{array}\right.
	$$
	As it follows from \cite[statement 2.9.1]{Kornejchuk1992} for all $p\in[1,\infty]$ the following inequality holds
	$$
	\|\tau_n(\cdot)\|_p
	\ge\|\cos n(\cdot)\|_p,
	$$
	and, consequently, for $p=\infty$
	\begin{equation}\label{1td14}
		\|\tau_n\|_C
		\ge1.
	\end{equation}
	
	From \eqref{1td13} and \eqref{1td14} we get
	\begin{equation}\label{1td15}
		\sqrt{a_n^2+b_n^2}=\frac{\left\|T_n\right\|_C}{\left\|\tau_n\right\|_C}\le\left\|T_n\right\|_C.
	\end{equation}
	
	Using the equations \eqref{1td9} and the estimates \eqref{1td12} and \eqref{1td15}  we have
	\begin{equation}\label{1td16}
		\left\|(T_n)_{\bar{\beta}}^\psi\right\|_2
		=\sqrt{\pi}	\left(\sum\limits_{k=1}^{n-1} \frac{a_k^2+b_k^2}{\psi^2(k)}+\frac{a_n^2+b_n^2}{\psi^2(n)}\right)^{\frac12}
		$$
		$$
		\le \sqrt{\pi}
		\left(2\sum\limits_{k=1}^{n-1} \frac{\left\|T_n\right\|^2_C}{\psi^2(k)}+\frac{\left\|T_n\right\|^2_C}{\psi^2(n)}\right)^{\frac12}=\sqrt{\pi}
		\left(\frac{1}{\psi^2(n)}+2\sum\limits_{k=1}^{n-1} \frac{1}{\psi^2(k)}\right)^{\frac12}\left\|T_n\right\|_C.
	\end{equation}
	
	Since the polynomials $T_n$  belongs to the ball $B_{2n+1}$ of the form \eqref{1td5}, from \eqref{1td16} it follows that
	\begin{equation}\label{1td17'}
		\left\|(T_n)_{\bar{\beta}}^\psi\right\|_2\le1.
	\end{equation}
	
	The embedding \eqref{1td6} is proved.
	
	The inequality \eqref{1td4} follows from the definition of the Berstein width $b_{2n}(C^\psi_{\bar\beta,2},C)$ and the embedding \eqref{1td6}. The relations \eqref{1td2}-\eqref{1td4} prove the inequalities \eqref{1t2}.

	To prove the asymptotic equations \eqref{1t4} under satisfying the condition \eqref{1t3} first of all we note  that
	\begin{equation}\label{1td17}
		\left(\sum\limits_{k=n}^\infty \psi^2(k)\right)^{\frac12}\le \psi(n)+\left(\sum\limits_{k=n+1}^\infty \psi^2(k)\right)^{\frac12}
	\end{equation}
	and
	\begin{equation}\label{1td18}
		\left(\frac1{\psi^2(n)}	+2\sum\limits_{k=1}^{n-1} \frac{1}{\psi^2(k)}\right)^{\frac12}\le \frac1{\psi(n)}+\left(2\sum\limits_{k=1}^{n-1} \frac{1}{\psi^2(k)}\right)^{\frac12}.
	\end{equation}
	
	From \eqref{1td18} we get
	\begin{equation}\label{1td19}
		\left(\frac1{\psi^2(n)}	+2\sum\limits_{k=1}^{n-1} \frac{1}{\psi^2(k)}\right)^{-\frac12}\ge \frac1{\frac1{\psi(n)}+\left(2\sum\limits_{k=1}^{n-1} \frac{1}{\psi^2(k)}\right)^{\frac12}}
		$$
		$$
		=\psi(n)-\left(\frac1{\frac1{\psi(n)}}-\frac1{\frac1{\psi(n)}+\left(2\sum\limits_{k=1}^{n-1} \frac{1}{\psi^2(k)}\right)^{\frac12}}\right)
		=
		\psi(n)-\frac{\left(2\sum\limits_{k=1}^{n-1} \frac{1}{\psi^2(k)}\right)^{\frac12}}{\frac1{\psi(n)}\left(\frac1{\psi(n)}+\left(2\sum\limits_{k=1}^{n-1} \frac{1}{\psi^2(k)}\right)^{\frac12}\right)}
		$$
		$$
		=\psi(n)\left(1-\frac{\left(2\sum\limits_{k=1}^{n-1} \frac{1}{\psi^2(k)}\right)^{\frac12}}{\frac1{\psi(n)}+\left(2\sum\limits_{k=1}^{n-1} \frac{1}{\psi^2(k)}\right)^{\frac12}}\right)
		\ge
		\psi(n)\left(1-\psi(n)\left(2\sum\limits_{k=1}^{n-1} \frac{1}{\psi^2(k)}\right)^{\frac12}\right).
	\end{equation}
	
	So, as it follows from \eqref{1t2} and \eqref{1td17}, on the one hand,
	\begin{equation}\label{1td20}
		P_{2n}(C^\psi_{\bar\beta,2},C)\le P_{2n-1}(C^\psi_{\bar\beta,2},C)
		\le
		\psi(n)\left(\frac1{\sqrt{\pi}}+\frac1{\sqrt{\pi}\psi(n)}\left(\sum\limits_{k=n+1}^\infty \psi^2(k)\right)^{\frac12}\right),
	\end{equation}
	and, on the other hand, by virtue of \eqref{1t2}
	\begin{equation}\label{1td21}
		P_{2n}(C^\psi_{\bar\beta,2},C)\ge
		\psi(n)\left(\frac1{\sqrt{\pi}}-\sqrt{\frac2{\pi}}\psi(n)\left(\sum\limits_{k=1}^{n-1} \frac{1}{\psi^2(k)}\right)^{\frac12}\right).
	\end{equation}
	The combination of \eqref{1td20} and \eqref{1td21} allows us to write equations
	\begin{equation}\label{1td22}
		P_{2n}(C^\psi_{\bar\beta,2},C)=\psi(n)\left(\frac1{\sqrt{\pi}}+\gamma^{(1)}_n\right),
	\end{equation}
	\begin{equation}\label{1td23}
		P_{2n-1}(C^\psi_{\bar\beta,2},C)=\psi(n)\left(\frac1{\sqrt{\pi}}+\gamma^{(2)}_n\right),
	\end{equation}
	in which for $\gamma^{(i)}_n, i=1,2,$ the following double inequalities hold
	\begin{equation}\label{1td24}
		-\sqrt{\frac2{\pi}}\psi(n)\left(\sum\limits_{k=1}^{n-1} \frac{1}{\psi^2(k)}\right)^{\frac12}\le\gamma^{(i)}_n
		\le
		\frac1{\sqrt{\pi}\psi(n)} \left(\sum\limits_{k=n+1}^\infty \psi^2(k)\right)^{\frac12}.
	\end{equation}
	
	If the condition \eqref{1t3} is satisfied, then by virtue of \eqref{1td22}-\eqref{1td24} the asymptotic equations \eqref{1t4} take place.
	
	Theorem 1 is proved.
\end{proof}

We note that the condition
\begin{equation}\label{25}
	\lim_{n\rightarrow\infty}\frac1{\psi^2(n)}{\sum\limits_{k=n+1}^\infty \psi^2(k)}=0
\end{equation}
is satisfied if $\psi(k)$  satisfies the condition $D_0$:
\begin{equation}\label{26}
	\lim_{k\rightarrow\infty}\frac{\psi(k+1)}{\psi(k)}=0.
\end{equation}

To make sure of this, let us put
\begin{equation*}\label{key}
	\varepsilon_n=\sup_{k\ge n}\frac{\psi(k+1)}{\psi(k)}.
\end{equation*}
By virtue of \eqref{26} $\varepsilon_n\downarrow0$ as $n\rightarrow\infty.$ So, we get
\begin{equation*}\label{key}
	\sum\limits_{k=n+1}^\infty \psi^2(k)=\psi^2(n)\left(\frac{\psi^2(n+1)}{\psi^2(n)}+\frac{\psi^2(n+2)}{\psi^2(n+1)}\frac{\psi^2(n+1)}{\psi^2(n)}+\ldots\right)
	$$
	$$
	\le
	\psi^2(n)\left(\varepsilon^2_n+\varepsilon^4_n+\ldots\right)
	=
	\psi^2(n)\frac{\varepsilon^2_n}{1-\varepsilon^2_n}=o(\psi^2(n)).
\end{equation*}

Let us show that for strictly decreasing sequences $\psi$ the fulfilment of condition $D_0$ of the form \eqref{26}  ensures the truth of the following equality
\begin{equation}\label{27}
	\lim_{n\rightarrow\infty} \psi^2(n) \sum\limits_{k=1}^{n-1} \frac1{\psi^2(k)} =0.
\end{equation}

To do this, we use Stoltz's theorem, according to which the relation \eqref{27} is followed from the following equality
\begin{equation}\label{28}
	\lim_{n\rightarrow\infty} \frac{\sum\limits_{k=1}^{n-1} \frac1{\psi^2(k)}-\sum\limits_{k=1}^{n-2} \frac1{\psi^2(k)}}{\frac1{\psi^2(n)}-\frac1{\psi^2(n-1)}}  =0.
\end{equation}

Since
\begin{equation*}\label{key}
	\frac{\sum\limits_{k=1}^{n-1} \frac1{\psi^2(k)}-\sum\limits_{k=1}^{n-2} \frac1{\psi^2(k)}}{\frac1{\psi^2(n)}-\frac1{\psi^2(n-1)}}
	=
	\frac{\frac1{\psi^2(n-1)}}{\frac1{\psi^2(n)}-\frac1{\psi^2(n-1)}}
	=
	\frac{\frac{\psi^2(n)}{\psi^2(n-1)}}{1-\frac{\psi^2(n)}{\psi^2(n-1)}},
\end{equation*}
then \eqref{28} follows from \eqref{26}.

In view of the above, we have the following statement.

\begin{corollary}
	Let $\bar{\beta}=\{\beta_k\}_{k=1}^\infty, \beta_k\in\mathbb{R},$ and the
	sequence $\psi(k)>0$ is strictly decreasing and satisfies the condition  $D_0$ of the form \eqref{26}. Then as $n\rightarrow\infty$ the asymptotic equalities \eqref{1t4} hold.
\end{corollary}

We give the corollaries of Theorem 1 in some important special cases.

\begin{theorem}\label{theorem2}
	Let $\bar{\beta}=\{\beta_k\}_{k=1}^\infty, \beta_k\in\mathbb{R},$ and $n\in\mathbb{N}.$ Then for all $\displaystyle r\ge\frac{n+1}2$  the following inequalities hold
	$$
	\frac1{\sqrt{\pi}}n^{-r}\left(
	1-\frac{4\left(1-\frac1n\right)^{2r}}{1+4\left(1-\frac1n\right)^{2r}}\right)^{\frac12}
	\le
	P_{2n}(W_{\bar\beta,2}^r,C)
	$$
	\begin{equation}\label{29}		\le
		P_{2n-1}(W_{\bar\beta,2}^r,C)
		\le
		\frac1{\sqrt{\pi}}n^{-r}\left(
		1+\frac{2+\frac1n}{\left(1+\frac1n\right)^{r}}\right)^{\frac12},
	\end{equation}
	where $P_N$ is any of the widths $b_N, d_N, \lambda_N$ or $\pi_N$.
\end{theorem}

\begin{proof}
	Let us put $\psi(k)=k^{-r}, r>1.$ Obviously, the condition \eqref{1t1} is satisfied. Since for $2r\ge n+1, n\in\mathbb{N},$
	$$
	\sum_{k=n+1}^{\infty}\frac1{k^{2r}}<\frac1{(n+1)^{2r}}+\int_{n+1}^{\infty}\frac{dt}{t^{2r}}=\frac1{(n+1)^{2r}}+\frac1{(2r-1)(n+1)^{2r-1}}
	=\frac1{(n+1)^{2r}}\frac{2r+n}{2r-1}
	$$
	\begin{equation}\label{29'}
		\le\frac1{(n+1)^{2r}}\frac{4r-1}{2r-1}
		\le\frac1{(n+1)^{2r}}\left(2+\frac1{2r-1}\right)
		\le
		\frac1{n^{2r}}\frac{2+\frac1{n}}{(1+\frac1n)^{2r}},
	\end{equation}
	then according to the right-hand side of the equality \eqref{1t2} of Theorem 1 we obtain an estimate
	\begin{equation}\label{30}
		P_{2n-1}(W_{\bar\beta,2}^r,C)\le
		\frac1{\sqrt{\pi}}\left(\sum_{k=n}^{\infty}\frac1{k^{2r}}\right)^{\frac12}
		\le 	
		\frac1{\sqrt{\pi}}n^{-r}\left(
		1+\frac{2+\frac1n}{\left(1+\frac1n\right)^{2r}}\right)^{\frac12}.
	\end{equation}
	
	On the other hand, for $\displaystyle r\ge\frac{n+1}2$ and $\psi(k)=k^{-r}$
	$$
	\frac1{\psi^2(n)} +2\sum\limits_{k=1}^{n-1} \frac1{\psi^2(k)}=
	n^{2r}+2\sum\limits_{k=1}^{n-1}k^{2r}
	\le n^{2r}+2\left((n-1)^{2r}+\int\limits_{1}^{n-1}t^{2r}dt\right)
	$$
	$$
	=
	n^{2r}+2\left((n-1)^{2r}+\frac{(n-1)^{2r+1}}{2r+1}-\frac1{2r+1}\right)
	<
	n^{2r}+2\left((n-1)^{2r}+\frac{(n-1)^{2r+1}}{n+2}\right)
	$$
	\begin{equation}\label{31}	
		<
		n^{2r}+4(n-1)^{2r}=n^{2r}\left(1+4\left(1-\frac1n\right)^{2r}\right).
	\end{equation}
	
	By virtue of the left part of the inequality \eqref{1t2} of Theorem 1 and the formula \eqref{31} we get an estimate
	\begin{equation}\label{32}
		P_{2n}(W_{\bar\beta,2}^r,C)
		\ge	
		\frac1{\sqrt{\pi}}n^{-r}\left(
		\frac{1}{1+4\left(1-\frac1n\right)^{2r}}\right)^{\frac12}
		=	\frac1{\sqrt{\pi}}n^{-r}\left(
		1-\frac{4\left(1-\frac1n\right)^{2r}}{1+4\left(1-\frac1n\right)^{2r}}\right)^{\frac12}.
	\end{equation}
	Combining the estimates \eqref{30} and \eqref{32} we obtain \eqref{29}.
	Theorem 2 is proved.
\end{proof}

Note that if a condition
\begin{equation}\label{33}
	\lim\limits_{n\rightarrow\infty}\frac rn=\infty
\end{equation}
is satisfied, then for $\psi(k)=k^{-r}$ the condition \eqref{26} is also satisfied because
\begin{equation*}\label{key}
	\frac{\psi(k+1)}{\psi(k)}=\left(\frac k{k+1}\right)^r=\left(1+\frac1k\right)^{-r}=\left(\left(1+\frac1k\right)^{k+1}\right)^{-\frac r{k+1}}\le e^{-\frac r{k+1}}\rightarrow0,\ k\rightarrow\infty.
\end{equation*}

Taking the limit, as $n\rightarrow\infty$, in the relations \eqref{29}, we obtain the following statement.

\begin{theorem}\label{theorem3}
	Let $\bar{\beta}=\{\beta_k\}_{k=1}^\infty, \beta_k\in\mathbb{R},$ $n\in\mathbb{N},$ and the condition \eqref{33} is satisfied.
	Then the following asymptotic equalities hold
	\begin{equation}\label{34}
		\left.\begin{array}{l}
			P_{2n}(W_{\bar\beta,2}^r,C)\\
			P_{2n-1}(W_{\bar\beta,2}^r,C)
		\end{array}\right\}=
		n^{-r}\left(\frac1{\sqrt{\pi}}+
		{\mathcal O}(1)\left(1+\frac1{n}\right)^{-r}\right),
	\end{equation}
	where $P_N$ is any of the widths $b_N, d_N, \lambda_N$ or $\pi_N$,  and ${\mathcal O}(1)$ are the quantities uniformly bounded in all parameters.
\end{theorem}

Note also that the equalities \eqref{34} are easy obtained from the formula \eqref{1t4} and estimates \eqref{29'} and \eqref{31}.

\begin{theorem}\label{theorem4}
	Let $\bar{\beta}=\{\beta_k\}_{k=1}^\infty, \beta_k\in\mathbb{R}$, $\alpha>0,$ $r>1,$ $n\in\mathbb{N}$ and be such that
	\begin{equation}\label{4t1}
		(n-1)^r>\frac1\alpha,
	\end{equation}
	then the following inequalities hold
	$$
	\frac1{\sqrt{\pi}}e^{{-}\alpha n^r}\left(
	1{-}\frac{2\gamma_{\alpha,r,n}e^{-2\alpha r (n-1)^{r-1}}}{1+2\gamma_{\alpha,r,n}e^{-2\alpha r (n-1)^{r-1}}}
	\right)^{\frac12}
	\le
	P_{2n}(C^{\alpha,r}_{\bar\beta,2},C)
	$$
	\begin{equation}\label{'29}
		\le
		P_{2n-1}(C^{\alpha,r}_{\bar\beta,2},C)
		\le	
		\frac1{\sqrt{\pi}}e^{-\alpha n^r}\left(1+e^{-2\alpha r n^{r-1}}\left(1+\frac1{2\alpha r n^{r-1}}\right)\right)^{\frac12},
	\end{equation}
	where $P_N$ is any of the widths $b_N, d_N, \lambda_N$ or $\pi_N$ and 
	\begin{equation}\label{4t4}
		\gamma_{\alpha,r,n}=\left(1+\frac{1}{\alpha r (n-1)^{r-1}}+e^{-2\alpha (n-1)^{r}}\max \left\{{e^{4\alpha}},\frac{e^2}{\alpha^{1{+}1/r}}\right\}\right).
	\end{equation}
\end{theorem}

\begin{proof}
	First of all, note that if $\alpha>0,$ $r>1,$ $n\in\mathbb{N}$ and satisfy the condition \eqref{4t1}, then for a quantity of the form
	\begin{equation}\label{p4t1}
		I_{n-1}:=\int\limits_{1}^{n-1}e^{2\alpha t^{r}}dt,\quad \alpha>0,\ r>1,
	\end{equation}
	the following inequality holds
	\begin{equation}\label{p4t2}
		I_{n-1}\le\frac{e^{2\alpha (n-1)^{r}}}{\alpha r (n-1)^{r-1}}+\max \left\{{e^{4\alpha}},\frac{e^2}{\alpha^{1{+}1/r}}\right\}.
	\end{equation}
	Indeed, integrating by parts we have
	\begin{equation}\label{p4t3}
		I_{n-1}=\frac1{2\alpha r}\int\limits_{1}^{n-1}  t^{1-r}de^{2\alpha t^{r}}=\frac1{2\alpha r}\left(\frac{e^{2\alpha (n-1)^{r}}}{(n-1)^{r-1}}-e^{2\alpha}\right)+\frac{r-1}{2\alpha r}\int\limits_{1}^{n-1}\frac{e^{2\alpha t^{r}}}{t^r}dt.
	\end{equation}
	For $0<2\alpha<1$, taking into account \eqref{4t1}, we obtain
	\begin{equation}\label{p4t4}
		\int\limits_{1}^{n-1}\frac{e^{2\alpha t^{r}}}{t^r}dt
		=
		\int\limits_{1}^{\alpha^{-1/r}}\frac{e^{2\alpha t^{r}}}{t^r}dt+\int\limits_{\alpha^{-1/r}}^{n-1}\frac{e^{2\alpha t^{r}}}{t^r}dt
		<\frac{e^2}{\alpha^{1/r}} +\alpha\int\limits_{\alpha^{-1/r}}^{n-1}e^{2\alpha t^{r}}dt\le \frac{e^2}{\alpha^{1/r}}+\alpha I_{n-1}.
	\end{equation}
	From \eqref{p4t3} and \eqref{p4t4} under condition  \eqref{4t1} 
	we get
	$$
	I_{n-1}<\frac{e^{2\alpha (n-1)^{r}}}{2\alpha r(n-1)^{r-1}}+\frac {e^2}{2\alpha^{1+1/r}}+\frac12I_{n-1}
	$$
	and
	\begin{equation}\label{p4t5}
		I_{n-1}<\frac{e^{2\alpha (n-1)^{r}}}{\alpha r(n-1)^{r-1}}+\frac {e^2}{\alpha^{1+1/r}},\quad 0<2\alpha<1.
	\end{equation}
	
	For $2\alpha\ge1$ we obtain
	$$
	\int\limits_1^{n-1}\frac{e^{2\alpha t^{r}}}{t^r}dt= \int\limits_{1}^{2^{1/r}}\frac{e^{2\alpha t^{r}}}{t^r}dt+\int\limits_{2^{1/r}}^{n-1}\frac{e^{2\alpha t^{r}}}{t^r}dt\le\int\limits_{1}^{2^{1/r}}{e^{2\alpha t^{r}}}dt+\frac12\int\limits_{2^{1/r}}^{n-1}{e^{2\alpha t^{r}}}dt=
	$$
	\begin{equation}\label{p4t6}
		=\frac12\int\limits_{1}^{2^{1/r}}{e^{2\alpha t^{r}}}dt+\frac12I_{n-1}<
		\frac{2^{1/r}-1}{2}e^{4\alpha}+\frac12I_{n-1}<\frac{e^{4\alpha}}2+\frac12I_{n-1}.
	\end{equation}
	From \eqref{p4t3} and \eqref{p4t6} under condition  \eqref{4t1} we have
	$$
	I_{n-1}<\frac{e^{2\alpha (n-1)^{r}}}{2\alpha r(n-1)^{r-1}}+\frac {e^{4\alpha}}{4\alpha}+\frac1{4\alpha}I_{n-1}<\frac{e^{2\alpha (n-1)^{r}}}{2\alpha r(n-1)^{r-1}}+\frac {e^{4\alpha}}{2}+\frac12I_{n-1}
	$$
	and
	\begin{equation}\label{p4t7}
		I_{n-1}<\frac{e^{2\alpha (n-1)^{r}}}{\alpha r(n-1)^{r-1}}+ {e^{4\alpha}},\quad 2\alpha\ge1.
	\end{equation}
	
	The inequality \eqref{p4t2} follows from \eqref{p4t5} and \eqref{p4t7}.
	
	For  $\psi(k)=e^{-\alpha k^{r}},\ \alpha>0, \ r>1, $ under condition  \eqref{4t1}, taking into account that for $r>1$ and $ n\in\mathbb{N}$
	\begin{equation*}\label{p4t8}
		\left(1+\frac 1n\right)^r-1>\frac {r}n,
	\end{equation*}
	we obtain 
	$$
	\psi^2(n)\sum\limits_{k=1}^{n-1}\frac1{\psi^2(k)}=
	e^{-2\alpha n^{r}}\sum\limits_{k=1}^{n-1}	e^{2\alpha k^{r}}
	\le e^{-\alpha n^{r}}\left( e^{2\alpha (n-1)^{r}}+\int\limits_{1}^{n-1}e^{2\alpha t^{r}}dt\right)
	$$$$
	\le e^{-2\alpha n^{r}} \left(e^{2\alpha (n-1)^{r}}\left(1+\frac{1}{\alpha r (n-1)^{r-1}}+e^{-2\alpha (n-1)^{r}}\max \left\{{e^{4\alpha}},\frac{e^2}{\alpha^{1{+}1/r}}\right\}\right)\right)
	$$
	\begin{equation}\label{p4t9}
		\le
		e^{-2\alpha (n-1)^{r}((1+\frac1{n-1})^r-1)} \gamma_{\alpha,r,n}
		\le
		\gamma_{\alpha,r,n}e^{-2\alpha r (n-1)^{r-1}}.
	\end{equation}
	
	Thus, by virtue of the left part of the inequality \eqref{1t2} of Theorem 1 and  \eqref{p4t9},  we obtain a required estimate from below for widths $P_{2n}(C^{\alpha,r}_{\bar\beta,2},C), \ \alpha>0, \ r>1,$   under condition  \eqref{4t1}
	$$
	P_{2n}(C^{\alpha,r}_{\bar\beta,2},C)\ge 
	\frac1{\sqrt{\pi}}\left(e^{2\alpha n^{r}}+2\sum\limits_{k=1}^{n-1}	e^{2\alpha k^{r}}\right)^{-\frac12}
	$$
	\begin{equation}\label{p4t10}
		\ge\frac1{\sqrt{\pi}}e^{{-}\alpha n^r}\left(
		1{-}\frac{2\gamma_{\alpha,r,n}e^{-2\alpha r (n-1)^{r-1}}}{1+2\gamma_{\alpha,r,n}e^{-2\alpha r (n-1)^{r-1}}}
		\right)^{\frac12}.
	\end{equation}
	
	As was shown in \cite[P. 163-164]{Stepanets1987}
	\begin{equation}\label{24}
		\sum_{k=n+1}^{\infty}e^{-\alpha k^r}<e^{-\alpha n^r}\left(1+\frac1{\alpha r n^{r-1}}\right)e^{-\alpha r n^{r-1}}, \quad r>1, \alpha>0, n\in\mathbb{N}.
	\end{equation}
	Therefore,
	\begin{equation}\label{p4t11}
		\frac1{\psi^2(n)}\sum\limits_{k=n+1}^\infty\psi^2(k)=e^{2\alpha n^r}\sum\limits_{k=n+1}^\infty e^{-2\alpha k^r}
		%\le	
		%e^{\alpha n^r}\sum_{k=n+1}^{\infty}e^{-\alpha k^r}
		<\left(1+\frac1{2\alpha r n^{r-1}}\right)e^{-2\alpha r n^{r-1}}.
	\end{equation}
	
	Thus,  by virtue of the right part of the inequality \eqref{1t2} of Theorem 1 and the formula \eqref{p4t11} we get an estimate
	\begin{equation}\label{p4t12}
		P_{2n}(W_{\bar\beta,2}^r,C)
		\le	
		\frac1{\sqrt{\pi}}e^{-\alpha n^r}\left(1+e^{-2\alpha r n^{r-1}}\left(1+\frac1{2\alpha r n^{r-1}}\right)\right)^{\frac12}.
	\end{equation}
	Combining the estimates \eqref{p4t10} and \eqref{p4t12} we obtain \eqref{'29}.
	Theorem 4 is proved.
\end{proof}

Taking the limit, as $n\rightarrow\infty$, in the relations \eqref{'29}, we obtain the following statement.

\begin{theorem}\label{theorem5}
	Let $\bar{\beta}=\{\beta_k\}_{k=1}^\infty, \beta_k\in\mathbb{R}$, $\alpha>0,$ $r>1,$ $n\in\mathbb{N}$ and the condition \eqref{4t1} is satisfied.  Then as $n\rightarrow\infty$ the following asymptotic equalities hold	
	\begin{equation}\label{5t1}
		\left.\begin{array}{l}
			P_{2n}(C^{\alpha,r}_{\bar\beta,2},C)\\
			P_{2n-1}(C^{\alpha,r}_{\bar\beta,2},C)
		\end{array}\right\}=
		e^{-\alpha n^r}\left(\frac1{\sqrt{\pi}}
		+ {\mathcal O}(1)\gamma_{\alpha,r,n}	e^{-\alpha r (n-1)^{r-1}}\right),
	\end{equation}
	where $P_N$ is any of the widths $b_N, d_N, \lambda_N$ or $\pi_N$ and $\gamma_{\alpha,r,n}$  is defined by \eqref{4t4} and ${\mathcal O}(1)$ are the quantities uniformly bounded in all parameters.
\end{theorem}

Note that the Theorem \ref{theorem5} complements the results of the works \cite{Serdyuk1999,Serdyuk_Bodenchuk2013JAT,Serdyuk_Sokolenko2011JAT,Shevaldin_1992,Stepanets_Serdyuk1995UMZh8},   which contain exact estimates for the widths of the classes of convolutions  with classical or generalized Poisson kernels. As it follows from the proofs of Theorems 4 and 5 the asymptotic equalities for widths in \eqref{5t1} are realized by trigonometric Fourier sums. The asymptotic equalities for deviations of Fourier sums on classes of generalized Poisson integrals $C^{\alpha,r}_{\beta,p}$ in the uniform metric are seen, for example, in \cite{Serdyuk_2005_8,Serdyuk_Stepanyuk2019,Stepanets2005} and others.

\bigskip

CONTACT INFORMATION

\medskip
A.S.~Serdyuk\\01024, Ukraine, Kyiv-4, 3, Tereschenkivska st.\\
serdyuk@imath.kiev.ua

\medskip
I.V.~Sokolenko\\01024, Ukraine, Kyiv-4, 3, Tereschenkivska st.\\
sokol@imath.kiev.ua
\end{document}